is (-11,45) a solution to the equation 3x+7y=-91? justify your answer
Answer by blwinbbbles(106)

You have a given x and y...so just plug them in and solve the equation
So:
3(-11) + 7(45) = -91
-33 + 315 = -91
282 = -91
that is not true..so these cordinates are not a solution to the equation